# Agentar-DeepFinance-300K: A Large-Scale Financial Dataset via Systematic Chain-of-Thought Synthesis Optimization

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have demonstrated remarkable general reasoning capabilities, holding significant potential for applications in the financial domain, a field that requires robust and reliable reasoning. It has been demonstrated that distilling high-quality chain-of-thought (CoT) rationales from advanced general reasoning models offers a promising and efficient path to the financial reasoning model. However, existing CoT synthesis methods suffer from shallow CoT sampling, leaving the question of how to construct a well-designed knowledge space for finance reasoning unexplored. In this paper, we present Agentar-DeepFinance-300K , a large-scale financial reasoning dataset characterized by its systematic CoT synthesis optimization. We first introduce a comprehensive CoT synthesis pipeline featuring Multi-perspective Knowledge Extraction (MKE) and Self-Corrective Rewriting (SCR) to generate exhaustive and deep financial reasoning trajectories. Furthermore, a systematic investigation, termed CoT Cube, is conducted to analyze critical factors that influence CoT effectiveness, such as necessity, length and synthesizer, yielding valuable insights for high-quality financial CoT construction. Experiments demonstrate that models trained on our Agentar-DeepFinance-300K achieve significant improvements on financial benchmarks. We publicly release Agentar-DeepFinance-300K , hoping to advance the research in financial reasoning models.
@@ 1 Introduction
Employing this method, we present Agentar-DeepFinance-300K , a large-scale financial dataset designed to push the boundary of financial reasoning models. Beyond the synthesizing CoTs, Agentar-DeepFinance-300K also incorporates multi-dimensional metadata annotation, such as complexity, quality and task type. Agentar-DeepFinance-300K is distinguished from existing financial reasoning datasets by its systematic CoT synthesis optimization. This includes building a comprehensive knowledge space for financial reasoning through our pipeline and refining CoT generation through the exploration of CoT Cube. Furthermore, to bridge the disparity between training data and real-world financial interactions, we additionally curated a dataset annotated by financial experts to reflect the financial capabilities required in real-world scenarios.
The remainder of this report is organized as follows. Section 2 provides an overview of our proposed Agentar-DeepFinance-300K . Section 3 details our comprehensive CoT synthesis pipeline. Section 4 presents an in-depth investigation into the factors influencing the effectiveness of CoT and the ablation study for the CoT synthesis pipeline. Section 5 discusses related works with this report. Section 6 summarizes the technical contributions and future works.
@@ 2 Dataset Overview
Figure 2: Illustration of the format of Agentar-DeepFinance-300K , which comprises three components: (1) the question, (2) the solution, including CoT and the final answer and (3) metadata, which encompasses multi-dimensional annotations such as complexity, quality and language.
@@ 3.1 Overview
Figure 3: Illustration of the pipeline for constructing Agentar-DeepFinance-300K .
@@ 6 Conclusion
This paper presents Agentar-DeepFinance-300K , a large-scale financial reasoning dataset constructed with a systematically optimized chain-of-thought (CoT) synthesis framework. We propose a pipeline incorporating multi-perspective knowledge extraction and self-corrective rewriting to produce comprehensive reasoning trajectories. Additionally, we systematically investigate critical factors influencing CoT effectiveness, providing empirical insights for high-quality financial CoT construction. Experiments demonstrate the effectiveness of our Agentar-DeepFinance-300K , with significant improvements on financial reasoning benchmarks of models trained on our dataset. Future work will focus on exploring multi-modal financial reasoning, incorporating visual financial content such as charts.

### v2

# Agentar-DeepFinance-100K: A Large-Scale Financial Dataset via Systematic Chain-of-Thought Synthesis Optimization

@@ Abstract
Recent advancements in large language models (LLMs) have demonstrated remarkable general reasoning capabilities, holding significant potential for applications in the financial domain, a field that requires robust and reliable reasoning. It has been demonstrated that distilling high-quality chain-of-thought (CoT) rationales from advanced general reasoning models offers a promising and efficient path to the financial reasoning model. However, existing CoT synthesis methods suffer from shallow CoT sampling, leaving the question of how to construct a well-designed knowledge space for finance reasoning unexplored. In this paper, we present Agentar-DeepFinance-100K , a large-scale financial reasoning dataset characterized by its systematic CoT synthesis optimization. We first introduce a comprehensive CoT synthesis pipeline featuring Multi-perspective Knowledge Extraction (MKE) and Self-Corrective Rewriting (SCR) to generate exhaustive and deep financial reasoning trajectories. Furthermore, a systematic investigation, termed CoT Cube, is conducted to analyze critical factors that influence CoT effectiveness, such as necessity, length and synthesizer, yielding valuable insights for high-quality financial CoT construction. Experiments demonstrate that models trained on our Agentar-DeepFinance-100K achieve significant improvements on financial benchmarks. We publicly release Agentar-DeepFinance-100K , hoping to advance the research in financial reasoning models. The project page is available at https://github.com/antgroup/Agentar-DeepFinance-100K.
@@ 1 Introduction
Employing this method, we present Agentar-DeepFinance-100K , a large-scale financial dataset designed to push the boundary of financial reasoning models. Beyond the synthesizing CoTs, Agentar-DeepFinance-100K also incorporates multi-dimensional metadata annotation, such as complexity, quality and task type. Agentar-DeepFinance-100K is distinguished from existing financial reasoning datasets by its systematic CoT synthesis optimization. This includes building a comprehensive knowledge space for financial reasoning through our pipeline and refining CoT generation through the exploration of CoT Cube. Furthermore, to bridge the disparity between training data and real-world financial interactions, we additionally curated a dataset annotated by financial experts to reflect the financial capabilities required in real-world scenarios.
The remainder of this report is organized as follows. Section 2 provides an overview of our proposed Agentar-DeepFinance-100K . Section 3 details our comprehensive CoT synthesis pipeline. Section 4 presents an in-depth investigation into the factors influencing the effectiveness of CoT and the ablation study for the CoT synthesis pipeline. Section 5 discusses related works with this report. Section 6 summarizes the technical contributions and future works.
@@ 2 Dataset Overview
Figure 2: Illustration of the format of Agentar-DeepFinance-100K , which comprises three components: (1) the question, (2) the solution, including CoT and the final answer and (3) metadata, which encompasses multi-dimensional annotations such as complexity, quality and language.
@@
Dataset $N_{\text{Q}}$ $N_{\text{R}}$ $N_{\text{A}}$Samples Proportion
Open-source Dataset
FinCorpus 86.25 1775.09 330.61 61,574 $53.64\%$
FinCUGE 99.17 498.44 232.40 18,527 $16.14\%$
Finance-Instruct-500K 151.49 920.85 558.15 10,343 $9.01\%$
FinQA 1051.05 1928.63 189.91 5,012 $4.37\%$
FinancialData 14.85 896.85 803.57 4,047 $3.53\%$
Quant-Trading-Instruct 65.97 2665.40 903.90 89 $0.08\%$
In-house Dataset
FinCAS 69.58 315.24 1.03 15,190 $13.23\%$
FinRA - - - - -
Table 1: Overview of Agentar-DeepFinance-100K data sources.
Data Source. The Agentar-DeepFinance-100K dataset is constructed from two primary sources: 99K samples from open-source data and 16K samples from our in-house data. For the open-source portion, extensive QA pairs were gathered from publicly available datasets, including FinCorpus [31], Finance-Instruct-500K [6], FinCUGE [17], FinQA [2], FinancialData [4] and Quant-Trading-Instruct [18]. These datasets underwent rigorous filtering to ensure quality and relevance to financial contexts, and instances overlapping with test sets in the benchmarks were systematically removed. The in-house data consists of two primary components: financial real annotations (FinRA) and content related to financial compliance and security (FinCAS). Table 1 summarizes the data sources, presenting some statistics such as average question length $N_{\text{Q}}$, reasoning trajectory length $N_{\text{R}}$ and answer length $N_{\text{A}}$. Reasoning trajectories were generated using QwQ-Plus.
Figure 3: Task composition.
Figure 4: Complexity distribution.
Task Diversity Analysis. Figure 4 presents the task composition within our Agentar-DeepFinance-100K , which encompasses six primary task domains: Knowledge QA, Natural Language Processing (NLP), Analysis & Interpretation, Math, Compliance & Security and Text Generation. The task distribution exhibits a significant concentration, with Knowledge QA and NLP collectively accounting for over $70\%$ of the dataset. In contrast, Text Generation tasks represent the smallest proportion, accounting for less than $1\%$. This distribution pattern aligns with observed user query tendencies in the financial domain, where interactions predominantly involve straightforward queries related to basic financial knowledge and information.
Distribution of Complexity. Figure 4 illustrates the distribution of complexity scores assigned to the Agentar-DeepFinance-100K , with scores ranging from 1 to 10. These complexity scores are generated by a large language model using a carefully crafted prompt designed to capture the relative complexity in solving the problems, rather than directly correlating with the difficulty of the tasks themselves. The results indicate that the complexity scores within the dataset are primarily concentrated in the range of 6 to 7, corresponding to medium-level questions. Additionally, a significant proportion of questions are scored between 3 and 5, which reflects lower levels of complexity. Furthermore, a smaller subset of the dataset includes questions with complexity scores exceeding 7, indicative of higher levels of complexity.
@@ 3.1 Overview
Figure 5: Illustration of the pipeline for constructing Agentar-DeepFinance-100K .
@@ 6 Conclusion
This paper presents Agentar-DeepFinance-100K , a large-scale financial reasoning dataset constructed with a systematically optimized chain-of-thought (CoT) synthesis framework. We propose a pipeline incorporating multi-perspective knowledge extraction and self-corrective rewriting to produce comprehensive reasoning trajectories. Additionally, we systematically investigate critical factors influencing CoT effectiveness, providing empirical insights for high-quality financial CoT construction. Experiments demonstrate the effectiveness of our Agentar-DeepFinance-100K , with significant improvements on financial reasoning benchmarks of models trained on our dataset. Future work will focus on exploring multi-modal financial reasoning, incorporating visual financial content such as charts.